Suppose investor A has $20,000 in his account and derives 4 utiles of Utility from this amount and would derive 5 Utiles of Utility if he had $40,000. He is faced with a choice to invest the $20,000 in a project that has 60% probability of earning a profit of $20,000 and 40% probability of losing $20,000?

Is the manager likely to invest in the project?

A project has expected risky cash flows of $45,000 in Perpetuity. Given that the risk adjusted rate of return for the project is 15%, and the risk free rate is 5%, What are the certainty equivalent cash flows in perpetuity?
1
Expert's answer
2020-03-27T10:31:26-0400

If the investor is faced with a choice to invest the $20000 in a project that has 60% probability of earning a profit of $20000 and 40% probability of loosing, then the value of the project is 20000*0.6 + (-20,000)*0.4 = $4,000, so the manager is likely to invest in the project.


A project has expected risky cash flows of $45,000 in perpetuity. Given that the risk adjusted rate of return for the project is 15%, and the risk free rate is 5%, the certainty equivalent cash flows in perpetuity are PV = 45,000/(0.15 - 0.05) = $450,000.